OpenStudy (anonymous):

What is 2 square root of 3 multiplied by the square root of 15 plus 3 square root of 5

OpenStudy (anonymous):

\[(2\sqrt{3}\times \sqrt{15})+3\sqrt{5}\] using order of operations...you must multiply first

OpenStudy (anonymous):

\[2\times \sqrt{3}\times \sqrt{15} = 2\times \sqrt{3\times15}\]

OpenStudy (anonymous):

\[2\times \sqrt{3\times 15}=2\times \sqrt{3\times 3\times 5}=2\times 3\times \sqrt{5}\]

OpenStudy (anonymous):

are you with me?